OpenStudy (anonymous):

Line passes through the points (1, 4) and (4, 9). Line m passes through the point (3, 2) and is parallel to line . What is the equation for line m?

OpenStudy (anonymous):

\[m=\frac{ y2-y1}{ x2-x1 }\] m=9-4 / 4-1 m=5/3 m=3 y-y1=m(x-x1) y-4=3(x-1) y-4=3x-3 y=3x+1 check it if you want.